-3x+2y=2 find the slope and y-intercept and graph
Answer by ewatrrr(24785)

HI,
*Hint: Solve for y to put the equation of the line into slope-intercept form
y = mx + b where m is the slope and b the y-intercept (value of y where the line crosses the y-axis)
.
-3x+2y=2
y = (3/2)x + 1
.
m = 3/2 and (0,1) a point on the line
As y increase by 3, x increases by 2.
(2,4) is also a point on the line